# Haptic bilateral teleoperation system for free-hand dental procedures

### v2

@@ Abstract
Free-hand dental procedures are typically repetitive, time-consuming and require high precision and manual dexterity. Robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload. However, robotic solutions for free-hand procedures remain limited or completely lacking, and their acceptance is still low. To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures (FH-HBTS). The system includes a dedicated mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site. By ensuring motion and force correspondence between the operator’s actions and the robot’s movements, monitored through visual feedback, we enhance the operator’s sensory awareness and motor accuracy. Furthermore, recognizing the need to ensure procedural safety, we limit interaction forces by scaling the motion references provided to the admittance controller based solely on measured contact forces. This ensures effective force limitation in all contact states without requiring prior knowledge of the environment. The proposed FH-HBTS is validated in a dental scaling procedure using a dental phantom. The results show that the system improves the naturalness, safety, and accuracy of teleoperation, highlighting its potential to enhance free-hand dental procedures.
Index Terms:
Haptics, teleoperation, interaction control, free-hand dental procedures, eye-hand coordination
@@ I Introduction
Traditionally, routine, prosthetic, periodontal, and some surgical dental procedures are performed free-hand. Routine procedures include the treatment of dental caries and endodontic therapy, characterized by the mechanical excavation of the damaged or infected dental structure, typically utilizing burs attached to a rotary dental handpiece, followed by the replacement of the lost structure with a synthetic filling material [2]. Prosthetic procedures, on the other hand, focus on the replacement of missing or severely damaged teeth through the use of prostheses, such as inlays, crowns, or bridges. These procedures involve the preparation of the tooth, which entails the removal of hard tissue from the damaged or healthy tooth using rotary instruments mounted on dental handpieces [3]. Periodontal procedures include scaling and root planing (SRP), which involve the removal of plaque and calculus from the surfaces of the teeth and the smoothing of root surfaces, typically employing manual hand instruments and inserts mounted on sonic or ultrasonic handpieces [4]. Finally, although robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement [5], many operators still prefer the free-hand method for dental implantology [6].
@@
Each of these procedures, even the simplest ones, require accuracy, extensively trained manual skills and dexterity.
@@
Therefore, the successful outcome of a procedure, as well as the comfort for the patient, depend on both human factors, i.e., the experience and skills of the human operator [7], the adequacy of the instrumentation used and the duration of the treatments [8], and environmental factors [9], as highlighted in Remark 1. In particular, the latter factors can significantly affect the operator’s comfort. In [10] a high prevalence of musculoskeletal disorders among dentists is reported. The main causes of these disorders can be traced back to an incorrect application of ergonomics, such as awkward postures, repetitive and prolonged movements, use of small or narrow tool handles, exertion of forces on hard or sharp objects, and exposure to vibrations [11]. Hence, when training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable posture [12]. In addition to this, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used.
@@
This is also apparent from emerging trends in teaching, which make use of virtual environments and haptic devices (HD)s for training new operators [13, 14].
Robots can achieve accuracy, precision, efficiency beyond human capabilities. For this reason, the role of robots is to improve dental procedures, both in terms of patient satisfaction, i.e. broader access to care, reduced treatment time, fast recovery, and in terms of operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, improved ergonomics [15], as shown in Figure 1.
@@ I-A Related works
Haptic teleoperations are a technology that enables the remote execution of complex tasks, employing a robotic system.
@@
By providing multiple types of sensory information, i.e. visual and haptic feedback, HBTSs enable the human operator to experience telepresence, i.e. the perception of interacting with the remote environment as if it were directly encountered [16]. Furthermore, by integrating different haptic feedback and guidance modalities [17, 18], and shared control approaches [19], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, while also improving operator comfort and confidence. For these reasons, such technology is applied in multiple domains, such as telesurgery, space exploration, and the handling of hazardous materials in nuclear or underwater environments. However, in the context of dentistry, the adoption of such systems is still very limited.
@@
The majority of research in robotics for dentistry is focused on oral implantology [20]. Robotic oral implant surgery offers significant benefits in terms of implant accuracy in partially [21] or totally edentulous patients, also compared to computer-assisted implant surgery (s-CAIS) systems [22]. Better placement accuracy, however, can be achieved at the expense of surgical efficiency, as shown in [23]. Such technologies also enable flapless dental implant placement, speeding up recovery time, reducing postoperative pain, and enhancing patient comfort [24]. [25] shows how a telerobotic-assisted drilling system can enhance the accuracy of the human operator while preserving awareness of the procedure through an effective transmission of cutting forces and vibrations.
@@
On the other hand, the state-of-the-art shows considerably less research effort in other free-hand procedures [26]. In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can also compensate for hand tremors and the lack of tactile feedback in laser ablation [28, 29], improving accuracy [30] and efficiency [31]. In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [32, 33], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [34]. In periodontal procedures, robotic systems have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied forces, as with other free-hand procedures. They can also improve the ergonomics for the operator, and comfort for the patient by speeding up recovery time. However, for such procedures, robotic technologies are still lacking [35], and there is no empirical evidence of such advantages.
@@
Thus, although robotics shows promise in dental practice, several limitations remain. First, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [36]. This is heavily influenced by the high costs of these robotic systems and the need to ensure their compatibility with existing clinical tools [37]. Second, the most widely used robots are fully autonomous, which require the development of state-of-the-art algorithms and advanced capabilities in perception, recognition, and exception handling [37]. In this context, the presence of a human is still necessary not only as a supervisor but also as an operator. In light of Remark 2 and Remark 3, and considering the impact on the operator acceptance [38], haptic teleoperations represent a viable solution. Moreover, the HBTSs would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [39] and enabling the effective and objective evaluation of operator skills during training [40]. Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of daily free-hand routine procedures and the less invasiveness, entailing better patient acceptance [41]. Considering these limitations, [26] recommends to shift the focus from oral implantology procedures towards less invasive ones, which are better suited as forerunners in the field, and also to invest in educational robotics research in academia as a valuable tool to overcome the barrier of acceptance of robotic systems among future dentists.
@@ I-B Contributions
Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and in light of Remark 3, in this paper we present a HBTS for free-hand dental procedures (FH-HBTS), including:
  • a dedicated mechanical end-effector, compatible with existing clinical tools and featuring an endoscopic camera, to improve visibility at the intervention site;
  • a virtualized force feedback, consistent with the interaction observed through the visual feedback of the eye-in-hand camera, which enhances the haptic experience for the operator;
  • a control system that ensures procedural safety, improves accuracy, and preserves manual dexterity, while providing motion correspondence between the operator’s movements and those of the end-effector, as observed in the camera view.
To the best of the authors’ knowledge, this is the only HBTS designed for free-hand dental procedures. Key contributions are:
  • a control system to ensure accurate control of the tool center point (TCP) and to guarantee motion correspondence, using a novel eye-hand coordination strategy;
  • a force limitation strategy that makes no previous assumption or estimation of the environment geometry, therefore accounting for the challenges outlined in Remark 1;
  • eye-hand force correspondence provided with saturation and filtering via a deadband function;
  • TRL-4 validation of the FH-HBTS on a free-hand procedure, consisting of dental scaling.
@@ II-A Notation
where ${\bf R}_{b}^{a}\in SO(3)$ is the rotation matrix from frame $a$ to frame $b$, and ${\bf p}_{b}^{a}\in\mathbb{R}^{3\times 1}$ represents $b$’s origin in frame $a$.
We let ${\bf x}_{b}^{a}$ be the pose of the frame $b$ with respect to the frame $a$, corresponding to ${\bf T}_{b}^{a}$, in the form:
${\bf x}_{b}^{a}=\begin{bmatrix}{\bf p}_{b}^{a}\\ {\bf q}_{b}^{a}\end{bmatrix},$ (2)
where ${\bf q}_{b}^{a}\in\mathbb{H}_{1}$ is the quaternion associated with ${\bf R}_{b}^{a}$.
@@
We denote by ${\bf v}^{a}_{b}\in\mathbb{R}^{6\times 1}$ the twist of the frame $b$ with respect to the frame $a$, in the form:
${\bf v}^{a}_{b}=\begin{bmatrix}\dot{\bf p}^{a}_{b}\\ {\boldsymbol{\omega}}^{a}_{b}\end{bmatrix},$ (3)
where $\dot{\bf p}^{a}_{b},{\boldsymbol{\omega}}^{a}_{b}$ are the linear and angular velocities, respectively.
@@ II-B Haptic device and robot reference frames
The frames used by the HD and the robot are reported in Figure 2. With regard to the HD, two frames are introduced:
  • $\mathcal{H}_{b}$ is the base reference frame, in which the HD’s stylus poses and the feedback forces are expressed;
  • $\mathcal{H}_{e}$ is the HD’s stylus frame.
Regarding the robot, four frames are defined:
  • $\mathcal{R}_{b}$ is the base reference frame, in which motion references are expressed;
  • $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the TCP;
  • $\mathcal{R}_{ed}$ is the desired robot’s end-effector frame, input to admittance control;
  • $\mathcal{R}_{c}$ is the compliant frame, output by the admittance control law.
@@ III Methodology
The FH-HBTS we propose, shown in Figure 3, is based on a position-force architecture [42], coupling the leader (LR) and follower (FR) robots. The LR is an impedance HD, sensing the pose of the human operator and rendering haptic feedback according to the remote interaction between the FR and the environment. The HD is controlled by a leader controller, featuring an eye-hand coordination controller (EHCC) and a haptic feedback controller (HFC). The former is responsible for capturing the HD’s stylus poses, filtering the human tremors, and generating the corresponding robot end-effector poses, so that the robot follows the stylus motion in coordination with the visual feedback. The latter is responsible for rendering the force feedback on the HD to let the human operator perceive the remote environment. Both controllers record data throughout the entire procedure. The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides the latter with a configurable degree of compliance on its target pose with respect to the environment. Such controller has three key features. First, it ensures the stability of the FR, even in case of loss of communication, since, in case of a lack of reference from the EHCC, the controller keeps following the last one received. Second, it avoids damage to the arm structure and manipulated environment, since it adapts the received references to ensure compliant interaction with the surroundings. Finally, it limits the interaction forces by scaling the received motion references through a free-hand force limitation strategy embedded in the controller itself.
With reference to Figure 3, depicting the complete system, the following sections examine its individual subcomponents in detail. In Section III-A and Section III-C, a detailed description of the objectives and operation of the EHCC and HFC controllers is provided. In Section III-B the robot interaction control is described.

### v3

@@ Abstract
Free-hand dental procedures are typically repetitive, time-consuming and require high precision and manual dexterity. Robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload. However, robotic solutions for free-hand procedures remain limited or completely lacking. To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures (FH-HBTS). The system includes a mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site. By ensuring motion and force correspondence between the operator’s and the robot’s actions, monitored through visual feedback, we enhance the operator’s sensory awareness and motor accuracy. Furthermore, to ensure procedural safety, we limit interaction forces by scaling the motion references provided to the admittance controller based solely on measured contact forces. This ensures effective force limitation in all contact states without requiring prior knowledge of the environment. The proposed FH-HBTS is validated both through a technical evaluation and an in-vitro pre-clinical study conducted on a dental model under clinically representative conditions. The results show that the system improves the naturalness, safety, and accuracy of teleoperation, highlighting its potential to enhance free-hand dental procedures.
@@ I Introduction
Traditionally, routine, prosthetic, periodontal, and some surgical dental procedures are performed free-hand. Routine procedures include the treatment of dental caries and endodontic therapy, while prosthetic procedures focus on the replacement of teeth through the use of prostheses. These procedures involve the preparation of the tooth, which entails the removal of hard tissue using rotary instruments mounted on dental handpieces [43]. Periodontal procedures include scaling and root planing (SRP), i.e., the removal of plaque and calculus from the teeth and the smoothing of root surfaces, typically employing hand instruments and inserts mounted on sonic or ultrasonic handpieces [53]. Finally, although robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement, many operators still prefer the free-hand method for dental implantology [2]. Each of these procedures requires accuracy, extensively trained manual skills, and dexterity.
@@
Therefore, the successful outcome of a procedure, as well as the comfort for the patient, depend on both human factors, i.e., the experience and skills of the human operator [29], the adequacy of the instrumentation used and the duration of the treatments [34], and environmental factors [54], as highlighted in Remark 1. In particular, the latter factors can significantly affect the operator’s comfort. In [55] a high prevalence of musculoskeletal disorders among dentists is reported. The main causes of these disorders can be traced back to an incorrect application of ergonomics, such as awkward postures, repetitive and prolonged movements, use of small or narrow tool handles, exertion of forces on hard or sharp objects, and exposure to vibrations [24]. Hence, when training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable postures [15]. In addition, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used.
@@
This is also apparent from emerging trends in teaching, which make use of virtual environments and haptic devices (HDs) for training new operators [37, 59].
In this context, the role of robots is to improve dental procedures, in terms of both patient satisfaction, i.e., broader access to care, reduced treatment time, and fast recovery, and operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, and improved ergonomics [1], as shown in Figure 1.
@@ I-A Related works
Haptic teleoperations are an enabling technology for the remote execution of complex tasks, employing a robotic system.
@@
By providing multiple types of sensory information, i.e. visual and haptic feedback, HBTSs enable the human operator to experience telepresence [45]. Furthermore, by integrating different haptic feedback and guidance modalities [67, 30], and shared control approaches [12], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, while also improving operator comfort and confidence. For these reasons, such technology is consolidated in, e.g., telesurgery, space exploration, and the handling of hazardous materials in nuclear or underwater environments. However, in the context of dentistry, its adoption is still very limited.
@@
The majority of research in robotics for dentistry is focused on oral implantology [6], mainly because robotic implant surgery offers significant benefits in terms of implant accuracy in partially [65] or totally edentulous patients. Such technology also enables flapless implant placement, accelerating recovery time, reducing postoperative pain, and enhancing patient comfort [58]. [28] shows that a telerobotic-assisted drill can enhance accuracy while preserving awareness through an effective transmission of cutting forces and vibrations.
On the other hand, the state-of-the-art shows considerably less research effort in other free-hand procedures, a trend also highlighted in existing reviews of dental robotics [3], which report a strong imbalance in favor of implantology. Despite this imbalance, the literature reports evidence of the potential benefits of robotic technologies in free-hand dental procedures. Such evidence is analyzed in the following and summarized in the supplementary materials [^1] This paper has supplementary downloadable material available at http://ieeexplore.ieee.org, provided by the authors. This includes a supplementary PDF format information file and a multimedia MOV format movie clip, showing the end-to-end system experimental demonstration. This material is 78.5 MB in size.. The analysis focuses on robotic systems explicitly developed for dental procedures and directly interacting with the oral environment, aiming at supporting or automating clinical free-hand procedures, while excluding works specifically addressing oral implantology. Only approaches supported by experimental, in vitro, or preliminary validation are considered, whereas purely virtual studies or robotic solutions not specifically tailored to dental procedures are not included.
@@
In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can compensate for hand tremors and the lack of tactile feedback in laser ablation [62, 63]. This can improve accuracy [35] and efficiency [68], and can enable automated tooth arrangement [70] and preparation for veneers and full crowns [40, 69]. In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [13, 14], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [26], and can support visual-guided tool navigation [23]. In periodontal procedures, robots have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied forces, as with other free-hand procedures. They can also improve the ergonomics for the operator, and comfort for the patient by reducing recovery time. However, for such procedures, robotic technologies are still lacking [33], and there is no empirical evidence of such advantages. Existing systems addressing force-regulated interaction with teeth are limited to adjacent tasks, such as tooth-brushing assistance [4], and are not designed for direct periodontal procedures.
@@
Thus, although robotics shows promise in dental practice, several limitations remain. First, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [61]. This is heavily influenced by the high costs of these systems and the need to ensure their compatibility with existing clinical tools [66]. Second, the most widespread robots are autonomous, which require the development of advanced capabilities in perception, recognition, and exception handling [66]. In this context, the presence of a human is still necessary not only as a supervisor but also as an operator. In light of Remark 2 and Remark 3, and considering the impact on the operator acceptance [32], haptic teleoperations represent a viable solution. Moreover, HBTSs would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [46] and enabling the effective and objective evaluation of operator skills during training [31]. However, as emerges from the literature analysis above, no robotic solutions based on haptic bilateral teleoperation have been reported for free-hand tasks. Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of daily free-hand routine procedures and the less invasiveness, entailing better patient acceptance [38]. Considering these limitations, [3] recommends to shift the focus from oral implantology towards less invasive procedures, which are better suited as forerunners in the field, and also to invest in educational robotics research as a valuable tool to overcome the barrier of acceptance of robotic systems among future dentists.
@@ I-B Contributions
Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and in light of the advantages introduced by haptic teleoperations, as in Remark 3, we propose a HBTS for free-hand dental procedures (FH-HBTS). The proposed system features a custom designed end-effector, compatible with existing clinical tools, and a dedicated haptic control system that ensures procedural safety, improves accuracy, and preserves manual dexterity. Therein, key contributions are: (1) an eye-hand coordination controller providing both motion and force correspondence between the HD’s workspace and the video stream coming from the eye-in-hand camera; with respect to existing HBTSs, such mapping is bilateral and drift-free, even after prolonged teleoperation sessions; (2) a force limitation strategy effectively reducing interaction forces; with respect to existing HBTSs, this strategy needs no previous assumption or online estimation of the environment geometry, therefore accounting for the challenges outlined in Remark 1; notably, this strategy is part of the onboard robot controller and is therefore compatible with passivity-based teleoperations; (3) an integrated TRL-5 system for free-hand dental procedures, validated on SRP, demonstrating technical feasibility and clinical suitability.
In light of the inclusion and exclusion criteria defined in Section I-A, and also reported in the supplementary materials, to the best of the authors’ knowledge, this is the first HBTS specifically complying with the special requirements of safety, accuracy, and dexterity characterizing free-hand dental procedures.
@@ II-A Notation
where ${\bf R}_{b}^{a}\in SO(3)$ is the rotation matrix from frame $a$ to frame $b$, with ${\bf R}_{a}^{b}=({\bf R}_{b}^{a})^{-1}\equiv({\bf R}^{a}_{b})^{\top}$, and ${\bf p}_{b}^{a}\in\mathbb{R}^{3\times 1}$ represents $b$’s origin in frame $a$.
We let ${\bf x}_{b}^{a}=\begin{bmatrix}({\bf p}_{b}^{a})^{\top},({\bf q}_{b}^{a})^{\top}\end{bmatrix}^{\top}$ be the pose of the frame $b$ with respect to the frame $a$, corresponding to ${\bf T}_{b}^{a}$, where ${\bf q}_{b}^{a}\in\mathbb{H}_{1}$ is the quaternion associated with ${\bf R}_{b}^{a}$.
@@
We denote by ${\bf v}^{a}_{b}=\begin{bmatrix}(\dot{\bf p}^{a}_{b})^{\top},({\boldsymbol{\omega}}^{a}_{b})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6\times 1}$ the twist of the frame $b$ with respect to the frame $a$, where $\dot{\bf p}^{a}_{b},{\boldsymbol{\omega}}^{a}_{b}\in\mathbb{R}^{3\times 1}$ are the linear and angular velocities, respectively.
We let ${\bf h}^{a}=\begin{bmatrix}({\bf f}^{a})^{\top},({\bf m}^{a})^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{6\times 1}$ be the wrench with respect to the frame $a$, where ${\bf f}^{a},{\bf m}^{a}\in\mathbb{R}^{3\times 1}$ are the forces and momenta, respectively.
${\bf I}_{n}$ indicates the $n\times n$ identity matrix.
@@ II-B Haptic device and robot reference frames
The frames used by the HD and the robot are reported in Figure 2. With regard to the HD, two frames are introduced: $\mathcal{H}_{b}$ is the base reference frame, in which the HD’s stylus poses and the feedback forces are expressed; $\mathcal{H}_{e}$ is the HD’s stylus frame. Regarding the robot, four frames are defined: $\mathcal{R}_{b}$ is the base reference frame, in which motion references are expressed; $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the tool center point (TCP); $\mathcal{R}_{ed}$ is the desired robot’s end-effector frame, input to admittance control; $\mathcal{R}_{c}$ is the compliant frame, output by the admittance control law.
@@ III Methodology
The FH-HBTS we propose, shown in Figure 3, is based on a position-position architecture [39], coupling the leader (LR) and follower (FR) robots by means of a spring–damper, which is a passive element. Therefore, in the absence of communication delays, this architecture has the advantage of being intrinsically stable [5, 25]. A formal proof of absolute stability can be found in [47]. In the proposed design, we assume negligible time delays and packet losses and, accordingly, do not integrate passivity-based control approaches [42]. This choice is motivated both by the validity of this assumption in healthcare facilities and by the conservative nature of passivity-based methods, which are known to worsen transparency and system usability in order to ensure stability under adverse network conditions [52]. Indeed, in precision free-hand dental procedures, this trade-off is generally undesirable, as it may impair fine force discrimination and the naturalness of the interaction. Nevertheless, the management of communication delays and packet losses represents an orthogonal direction to the one investigated in this work. Indeed, under non-ideal network conditions, the proposed architecture can be readily extended to enforce time-domain passivity of the communication channel, for instance by adopting a TDPA-based approach [7], as demonstrated in Section IV-B.
The LR is an impedance HD, sensing the pose of the operator and rendering haptic feedback according to the remote interaction between the FR and the environment. The HD is controlled by a leader controller, featuring an eye-hand coordination controller (EHCC) and a haptic feedback controller (HFC). The former is responsible for capturing the HD’s stylus poses, filtering human tremors, and generating the corresponding robot end-effector poses, so that the robot follows the stylus motion in coordination with the visual feedback. The latter is responsible for rendering the force feedback on the HD to let the human operator perceive the remote environment. Both controllers record data throughout the entire procedure. The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides a configurable degree of compliance on its target pose with respect to the environment and improves safety through dedicated mechanisms and emergency protocols.
With reference to Figure 3, depicting the complete system, the following sections examine its individual subcomponents in detail. Section III-A and Section III-C provide a detailed description of the objectives and operation of the EHCC and HFC controllers. Section III-B describes the robot interaction control. Finally, Section III-D presents the design of the mechanical end-effector for the FR.